OpenStudy (anonymous):

Solve this system of equations. x+y+z=-1 2x+5y+2z=1 -x+6y-3z=0

myininaya (myininaya):

1 1 1 | -1 2 5 2 | 1 -1 6 -3 | 0 = 1 1 1 | -1 here i multiplied row 1 by -2 and added to row 2 0 3 0 | 3 -1 6 -3 | 0 = 1 1 1 | -1 here i added row 1 to row 3 0 3 0 | 3 0 7 -2 | -1

myininaya (myininaya):

= 1 1 1 |-1 0 1 0 | 1 divided row 2 by 3 0 7 -2 | -1 = 1 1 1 | -1 0 1 0 | 1 here multiplied row 2 by -7 and added to row 3 0 0 -2| -8

myininaya (myininaya):

1 1 1 | -1 0 1 0| 1 0 0 1| 4 divided row 3 by -2 = 1 0 1|-2 0 1 0| 1 here multiplied row 2 by -1 and added to row 1 0 0 1|4

myininaya (myininaya):

= 1 0 0|-6 0 1 0| 1 here mutlpied row 3 by -1 and added to row 1 0 0 1|4

myininaya (myininaya):

first column is the x column second column is the y column third column is the z column

myininaya (myininaya):

so you have x=-6, y=1 ,z=4

myininaya (myininaya):

you should check my arithmetic though

OpenStudy (anonymous):

ok. thank you!